# Gated InAs quantum dots embedded in surface acoustic wave cavities for low-noise optomechanics

### v1

@@
In conclusion, we have successfully incorporated low-noise gated quantum dots into high-quality-factor surface acoustic wave resonators to create fully integrated optomechanical devices, aimed at microwave-to-optical transduction. We showed that the conductive layers required for QD gating negligibly impact the SAW resonator performance. Specifically, with suitable choice of the location and the doping concentration of the doped layer, SAW generation and attenuation are virtually unaffected compared to reference systems without the QD gate layers. As a result, acoustic resonators retain excellent quality factors. Further, the additional fabrication processes required for the SAW cavities do not negatively impact the QD performance; QDs retain near transform-limited linewidths. This is critically important when operating the system as a microwave-to-optical transducer which requires illuminating the QD with a strong red-detuned pump; a narrow QD linewidth substantially reduces the rate at which this red-detuned pump directly excites the QD without phonon involvement. Therefore, the near transform-limited QDs improve the rate of getting quantum information converted to optical photons from acoustic phonons, compared to directly exciting the QD. This is a key step for building efficient low-noise quantum transducers with QDs and SAWs.
[…]
Funding National Research Council; National Institute of Standards and Technology.
[…]
@@
Acknowledgments The authors thank Adam McCaughan for providing the Python layout design code PHIDL. The authors also thank John Teufel for the discussion on microwave transmission lines.
[…]
@@
Data availability Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
[…]
Supplemental document See Supplement 1 for supporting content.
Gated InAs quantum dots embedded in surface acoustic wave cavities for low-noise optomechanics: supplemental document
1 Sample Structure and Fabrication Process
@@
The MBE-grown wafer is then processed for the combined gated QD and SAW device. First, the SAW DBR (mirror) trenches are defined by electron beam lithography (EBL) and etched by reactive-ion etching (RIE). Then, ohmic contacts to the n-doped GaAs layer are made by etching vias around the SAW resonators down to approximately 100 nm above the GaAs:Si layer, and then depositing approximately 400 nm of Ni/AuGe/Ni/Au by electron-beam evaporation. The metals are thermally annealed in forming gas (95% N2 and 5% H2) at 420 C for 1 minute for optimum contact resistance. To finish the Schottky diode structure, a thin layer of Ti/Au (3 nm/7 nm) is evaporated inside the acoustic resonator region, along with a thick layer of Ti/Au (10 nm/250 nm) outside the resonator region to connect to large wirebond pads. In the next step, the n-doped layer is etched away everywhere outside the SAW resonator region so that the on-chip microwave coplanar waveguides (CPWs) and wirebond pads are unaffected by the doped layer. The CPW traces are then defined by photo-lithography and lifted-off with 450nm of Ti/Au. The thickness of the CPW metal is chosen to be thicker than the etched depth so that it is connected across the step. Finally, 20 nm of either Au or Al are deposited to fabricate the IDTs. The IDT structure is defined by EBL and then formed using a lift off process. We used Au IDTs for measurements shown in Figs. 2a-c and Al IDTs for measurements shown in all other figures.
2 Theoretical estimates for SAW propagation losses and electromechanical coupling in our system.

### v2

@@ Discussion
In conclusion, we have successfully incorporated low-noise gated quantum dots into high-quality-factor surface acoustic wave resonators to create fully integrated optomechanical devices, aimed at microwave-to-optical transduction. We showed that the conductive layers required for QD gating negligibly impact the SAW resonator performance. Specifically, with suitable choice of the location and the doping concentration of the doped layer, SAW generation and attenuation are virtually unaffected compared to reference systems without the QD gate layers. As a result, acoustic resonators retain excellent quality factors. Further, the additional fabrication processes required for the SAW cavities do not negatively impact the QD performance; QDs retain near transform-limited linewidths. This is critically important when operating the system as a microwave-to-optical transducer which requires illuminating the QD with a strong red-detuned pump; a narrow QD linewidth substantially reduces the rate at which this red-detuned pump directly excites the QD without phonon involvement. Therefore, the near transform-limited QDs improve the rate of getting quantum information converted to optical photons from acoustic phonons, compared to directly exciting the QD. This is a key step toward truly leveraging the excellent properties of semiconductor QDs for optomechanics.
Acknowledgments
@@
The authors thank Adam McCaughan for providing the Python layout design code PHIDL. The authors also thank John Teufel for the discussion on microwave transmission lines.
Data availability
@@
Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
Supplementary Information
I Sample Structure and Fabrication Process
@@
The substrate used in this work is grown by molecular beam epitaxy (MBE), with layer structures described in supplemental information. The MBE-grown wafer is then processed for the combined gated QD and SAW device. First, the SAW DBR (mirror) trenches are defined by electron beam lithography (EBL) and etched by reactive-ion etching (RIE). Then, ohmic contacts to the n-doped GaAs layer are made by etching vias around the SAW resonators down to approximately 100 nm above the GaAs:Si layer, and then depositing approximately 400 nm of Ni/AuGe/Ni/Au by electron-beam evaporation. The metals are thermally annealed in forming gas (95% Ar and 5% H2) at 420 C for 1 minute for optimum contact resistance. To finish the Schottky diode structure, a thin layer of Ti/Au (3 nm/7 nm) is evaporated inside the acoustic resonator region, along with a thick layer of Ti/Au (10 nm/250 nm) outside the resonator region to connect to large wirebond pads. In the next step, the n-doped layer is etched away everywhere outside the SAW resonator region so that the on-chip microwave coplanar waveguides (CPWs) and wirebond pads are unaffected by the doped layer. The CPW traces are then defined by photo-lithography and lifted-off with 450nm of Ti/Au. The thickness of the CPW metal is chosen to be thicker than the etched depth so that it is connected across the step. Finally, 20 nm of either Au or Al are deposited to fabricate the IDTs. The IDT structure is defined by EBL and then formed using a lift off process. We used Au IDTs for measurements shown in Fig. 2a-c and Al IDTs for measurements shown in all other figures.
II Theoretical estimates for SAW propagation losses and electromechanical coupling in our system.
@@
III Microwave Reflection Model
The microwave reflection and transmission measurements shown in Fig. 2 are conducted with commercial vector network analyzer. The internal and external quality factor can then be extracted by fitting the reflection spectra according to Eqn. 1 in Ref. [55]:
$S_{11}(f)=\frac{(Q_{e}-Q_{i})/Q_{e}+2iQ_{i}(f-f_{0})/f}{(Q_{e}+Q_{i})/Q_{e}+2iQ_{i}(f-f_{0})/f}$
The size of the dip in $|S_{11}|^{2}$ reflection spectrum provides an upper bound on the electromechanical coupling efficiency, which can be improved by impedance matching with tunable superconducting microwave cavity [62].
IV Optical Measurement Setup
The optical measurements shown in Fig. 3 are conducted with conventional confocal microscope setup, with details described in Ref. [21]. The microwave power used for measurements in Fig. 3c) and 3d) are -29 dBm and -26 dBm, respectively. For the cavity mode studied in Fig 3d), microwave reflection measurement shows a $Q_{i}$ of 18,396 and $Q_{e}$ of 548,077, corresponding to a total Q of approximately 17,798. With methods described in Ref. [46] appendix D, We estimate the single-phonon coupling rate $g_{0}$ to be approximately 42 kHz.